1/5ln(x)-2(ln(x+2)) as a single logarithm. PLEASE HELP! and please put steps. I am so lost.
Answer by lwsshak3(11628)

Write as a single logarithm.
1/5ln(x)-2(ln(x+2))
1/5ln(x)=ln(x^(1/5))
2(ln(x+2))=ln(x+2)^2
subtraction in logs means division in operations
Placing under a single log:
1/5ln(x)-2(ln(x+2))=ln[x^(1/5)/(x+2)^2]